OpenStudy (anonymous):

A company will need $60,000.00 in 5 years for a new addition. To meet this goal, the company deposits money in an account today that pays 3% annual interest compounded quarterly. Find the amount that should be invested to total $60,000.00 in 5 years

OpenStudy (anonymous):

the formula for compound interest (if i remember correctly) is: Total = Invested Amount * (1 + annual_rate/100)^(time) since it's compounded quarterly for 5 years, then it's compounded 20 times, so the total, if your investment is P, will be P * (1 + 3/100)^(20) = P * (1.03)^20 = 60,000 Hence P = 60,000/(1.03)^20 i could be wrong though